OpenStudy (anonymous):

Giving medal or best answer Solve 6/n^2 - 1/n = -1/3n

OpenStudy (anonymous):

Try combining terms on the left hand side of the equation first \[\frac{ 6 }{ n ^{2} } - \frac{ 1 }{ n } = -\frac{ 1 }{ 3n }\] what do you get when you do that?

OpenStudy (anonymous):

(6-n)/(n^2)=-1/(3n) (6-n)=(-1/3)n 6=n-n/3 6=2n/3 n=9